OpenStudy (anonymous):

Use the formulas a = x^2 – y^2, b = 2xy, c = x^2 + y^2 to answer this question. One leg of a right triangle has length 11 and all sides are whole numbers. Find the lengths of the other two sides.

jimthompson5910 (jim_thompson5910):

Notice how 6^2 - 5^2 = 36 - 25 = 11 so we see that x = 6 and y = 5. Use these values to find b and c

OpenStudy (anonymous):

I dont understand

jimthompson5910 (jim_thompson5910):

the formulas a = x^2 – y^2, b = 2xy, c = x^2 + y^2 are used to find a,b,c which are sides of the right triangle c is the longest side (hypotenuse)

jimthompson5910 (jim_thompson5910):

a = 11 ---> x^2 - y^2 = 11 doing a bit of trial and error, you can find that x = 6 and y = 5 make the equation x^2 - y^2 = 11 true